Status Report

NASA Space Station On-Orbit Status 3 February 2005

By SpaceRef Editor
February 3, 2005
Filed under , , ,
NASA Space Station On-Orbit Status 3 February 2005

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

FE Salizhan Sharipov conducted the second experiment session with the Russian/German TEKh-20 Plasma Crystal-3 (PK-3) science payload. [First thing in the morning, Salizhan again activated the PK-3/N turbopump in the Service Module (SM) Transfer Compartment (PkhO), which keeps a vacuum inside the work chamber (ZB) in the SM RO (Work Compartment). Then he set up the experiment, supported by tagup with the ground. Video recording began at experiment initiation plus ~6 min. The turbopump will be deactivated tonight at ~4:25pm EST. The experiment, which is run in automatic mode, is performed on plasma, i.e., fine particles charged and excited by HF (high frequency) radio power inside the evacuated work chamber. Main objective today was to study dispersion properties of the plasma dust cloud at various pressures and particle quantities with superimposition of an LF (low frequency) harmonic electrical field.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

Sharipov configured the Progress M-51 (16P) “Rodnik” tank system for liquid waste transfer from the SM holding tank to the empty Rodnik water tanks. The transfer, from seven EDV-U urine containers to SM tankage and thence to the Progress, proceeded during the day, driven by a compressor. An additional eight filled EDV-Us are slated for 16P disposal with contents intact. [The collapsed bladders of the Rodnik BV1 & BV2 water storage tanks in the Progress were pressurized yesterday as a longer-term leak-checkup preparatory to the liquid waste transfer to the tanks for disposal. Each of the two spherical Rodnik tanks consists of a hard shell with a soft membrane (bladder) composed of elastic fluoroplastic. The bladder is used to expel water from the tank by compressed air pumped into the tank volume surrounding the membrane.]

CDR Leroy Chiao’s main activity today was the uploading of new firmware to the receiver/processor of SIGI-1 (Space Integrated GPS/Inertial Navigation System), as was done for SIGI-2 on 11/11/04. [The activity was performed in five separate steps, viz.: relocation of OpsLAN to gain access to the Lab’s starboard Zone 7 and rack LAB1S7, rotation of the rack for accessing a cable assembly, loading the firmware via the cabling, restowing the cable assembly and rotating the rack back, and finally relocating OpsLAN to its former location on rack LAB1S6. The GPS upload consists of the new Honeywell OFP (operational flight program) firmware plus two additional modules (“Trimble” card firmware & GSFC attitude firmware). After SIGI-1 power-up, error indications appeared and are now being investigated by specialists.]

Leroy replaced the microdrive and changed the lens on the EarthKAM (Earth Knowledge Acquired by Middle school students, EK) system’s Kodak DSC 760 digital still camera at the Lab window, going from 50mm to the 180mm-lens configuration. Besides using the DSC 760 instead of the earlier Kodak ESC 460C electronic still camera, EarthKAM has switched to an A31p NGL (next generation laptop) with new software. [About 160 schools worldwide are currently using EK. Over 900 images have been requested so far. Downlinked images are posted on the EK website ( ). EarthKAM deactivation and stow is scheduled for 2/5 (Saturday).]

The FE updated the RODF (Russian operations data file) book on Fire Protection System (SPPZ) by updating the procedure for donning Russian IPK-1M gas masks.

Salizhan also performed the periodic (weekly) replenishing of the Elektron system’s water supply for electrolysis, his 13th, filling the KOV thermal loops EDV container today with water from SM Rodnik tank #1 (BV1). [The procedure is specially designed to prevent air bubbles from getting into the BZh liquid unit where their pressure spikes, from collapsing, could cause micropump impeller cavitation and Elektron shutdown, as happened numerous times in the past. Usually, when the Rodnik’s potable water is not used, the EDV water is condensate drawn from the BKO multifiltration/purification column and the air/liquid separator unit (GZhS) while the crewmember checks for any air bubbles in the EDV (and, if visible, estimates their number). Elektron water is also supplied from USOS condensate in a CWC (collapsible water container) that is checked for its contents of air bubbles and is rejected if the estimated total air bubble volume is more than 30 cubic centimeters (1 cm air bubble is about 0.5 ccm).]

Leroy conducted the weekly data take with the two newly delivered CSA-CP units (Compound Specific Analyzer-Combustion Products, #1018 & #1019), with the purpose to monitor the ongoing decontamination (outgassing) of the deployed instruments. [CSA-CP measures O2 (oxygen), CO (carbon monoxide), HCN (hydrogen cyanide), and HCl (hydrogen chloride).]

The CDR also completed his daily check of the Total Dose reading and End File values of the TEPC (Tissue Equivalent Proportional Counter), which he relocated to the Node on 1/31, and called the data down at the evening DPC (daily planning conference). [This is currently a daily requirement since the UOP (utility outlet panel) near the TEPC’s temporary location (for two weeks) does not allow automated telemetry monitoring by the ground.]

Salizhan held his weekly IMS (inventory management system) tagup with ground specialists, discussing open issues concerning identification of equipment and storage locations for IMS updating. [Today’s topics dealt with such items as barcode identification and location of the EVA-12 film with views of the exhaust vent residues (no digital images were taken of the latter), location of IELKs (individual equipment & liner kits, Russian: USIL) location, quantity of food waste bags, etc.]

Chiao continued cargo prepacking for the LF-1 Shuttle mission (STS-114), the first RTF (return-to-flight) checkout flight. The activity is supported by an updated prepack list uplinked to the crew overnight. [Return equipment is placed in designated bags and predominantly staged on the FGB floor and in one of the ZSRs (zero-G storage racks) in the Lab (LAB1P4), which is slated to return on LF-1.]

The FE completed the routine SOZh/ECLSS maintenance in the SM, including the ASU toilet facilities. Later, he also prepared the regular IMS “delta” file for the daily automated export/import to the three IMS databases on the ground.

The CDR/SO filled out the regular weekly FFQ (food frequency questionnaire), his 14th, which keeps a log of his nutritional intake over time on special MEC (medical equipment computer) software. [The FFQ records amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. IBMP-recommended average daily caloric value of the crew’s food intake is 2200-2300 cal. If larger quantities of juices and fruits are taken into account, the value can go to 2400-2500 cal.]

Salizhan transferred the accumulated data files of his BIO-5 Rasteniya-2 (“Plants-2”) plant growth experiment to the Packet laptop for subsequent downlink to the ground via Regul comm.

The crew performed their daily 2.5-hr. physical exercise program on TVIS treadmill (aerobic), RED exerciser (anaerobic) and VELO cycle with bungee cord load trainer (both aerobic and anaerobic). Salizhan’s daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 4 of a new set) and one hour on VELO.

Chiao then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.

As reported on 12/14/04, the U.S. Airlock (A/L) is off limits for spacewalks at this time due to rust corrosion in the heat exchanger (HX) between the Lab’s high-pressure ITCS (internal thermal control system) and the lower pressure cooling loop of an EMU spacesuit when it is connected in the A/L. A new and improved HX has now been officially manifested on Progress M-52 (17P), due to launch in three weeks. [Chiao and Sharipov are qualified to perform the HX installation, as will be the Expedition 11 crew of Krikalev and Phillips. Afterwards, the A/L will have to be re-certified prior to LF-1 (STS-114), or else the LF-1 EVAs will have to be conducted from the Shuttle airlock.]

The Elektron is functioning nominally. [The electrolysis machine operates in 32-amp mode until 2/9, when the next regeneration cycle of the Micropurification Unit (BMP) starts, which uses the same vacuum vent valve. Following BMP filter bed regeneration, the Elektron remain off. On 2/15, the station interior will be repressurized with O2 from Progress 16 by ~10mmHg/Torr. Ten days later (2/25), a second O2 repress of ~15mmHg will be performed, depleting the remaining 16P oxygen. Progress undocking is currently scheduled for 2/26. The Elektron O2 generator will then be re-activated in early March.]

Energia/TsUP testing of the SM propulsion system (ODU) manifold 1 (both fuel and oxidizer) continued today. [The test is intended to characterize the manifold’s ability to hold a pressure over a period of time. No pressure decay has been observed to date, indicating zero leaks. The test will run for approximately four days, followed by an identical test on manifold 2, scheduled to begin on or about 2/7. This could become a routine activity as determined by the system specialists.]

The station continues to fly “sideways” in earth-fixed LVLH YVV attitude (local vertical local horizontal/y-axis in velocity vector) until 2/8 (Tuesday) when it maneuvers to sun-oriented XPOP (x-axis perpendicular to orbit plane) as the solar Beta angle dips below -60 deg magnitude.

Volume I (402 pages) of “Rockets and People” (Rakety i lyudi), the remarkable four-volume memoirs of 93-year old Russian space pioneer Academician Boris E. Chertok, translated from the original Russian, has now been published in the NASA History Series with the support of the ISS Program Office. [Contact: NASA Center for AeroSpace Information, 7121 Standard Drive, Hanover, MD 21076, 301-621-0390, , Online Order Form: , NASA Report #: NASASP20054110. Title: Rockets and People. The price code is A03 (within U.S. $27.50 plus $2.00 shipping & handling; outside U.S. $55.00 plus $17.00 S&H). This book also may be purchased from the NASA Information Center, Code CMI-1, NASA Headquarters, 300 E Street SW, Room 1H23, Washington, DC 20546-0001, (202) 358-0000. Order NASA SP-2005-4110.]

Today’s CEO (crew earth observations) photo targets, no longer limited in the current LVLH attitude, were Glacial features – South Libya (this portion of Libya was glaciated approximately 450 million years ago. Meltwater cut drainage channels beneath the ice sheet that were subsequently filled in by later deposits. Still later erosion of the region has exposed these fossil river channels. Nadir photography of the area will be useful for detection of these channels, which are expressed as subtle sinuous features on the landscape), Northern Sierra Nevada, CA (Dynamic Event. This overpass provided an opportunity for photography of the northern peaks of the Sierra. The western U.S. has received large amounts of snowfall during this winter, and imagery of the snowpack will be useful for assessment of the fresh water supply for this coming year), and Stardust Landing Site, Utah (looking to the left of track for the Stardust landing ellipse area. Mapping swaths of the region are useful for establishing geographic context and landmark positions for surface water mapping).

CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 10 crew visit:

Expedition 10 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Upcoming Key Events:

  • Progress M-51 (16P) undocking & destructive reentry — 2/26/05;
  • Progress M-52 (17P) launch — 2/28/05.
  • EVA-13 — 3/25/05;
  • Soyuz TMA-6 (10S) launch — 4/15/05 with Expedition 11 (CDR Sergei Krikalev, FE/SO John Phillips);
  • Soyuz TMA-5 (9S) undock — 4/25/05 with Exp. 10 crew (after 193 days on orbit, 191 days on board ISS);
  • Progress M-53 (18P) launch — 6/10/05;
  • Progress M-54 (19P) launch — 8/24/05;
  • Soyuz TMA-7 (11S) launch — 9/27/05.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit (as of this morning, 6:50am EST [= epoch]):

  • Mean altitude — 357.3 km
  • Apogee height — 363.9 km
  • Perigee height — 350.8 km
  • Period — 91.69 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0009686
  • Solar Beta Angle — -67.9 deg (magnitude decreasing)
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 56 m
  • Revolutions since FGB/Zarya launch (Nov. 98) — 35466

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.